You own a portfolio that has a total value of $185,000 and it is invested in Stock D with a beta of .91 and Stock E with a beta of 1.33. The beta of your portfolio is equal to the market beta. What is the dollar amount of your investment in Stock D?

Multiple Choice

  • $29,732.37

  • $34,687.50

  • $145,357.14

  • $39,642.86

  • $19,821.88
Answer #1

Let investment in D=$x

Hence investment in E=(185,000-x)

Portfolio beta=Respective beta*Respective investment weight

1=(x/185,000*0.91)+(185,000-x)/185,000*1.33[Beta of market=1;Beta of risk-free assets=0]

(1*185000)=0.91x+246050-1.33x

185,000=0.91x+246050-1.33x

x=(246050-185,000)/(1.33-0.91)

=$145,357.14=investment in D(Approx).